OpenStudy (anonymous):

Find and solve a system of linear equations to answer the question below. Joanne buys one kilogram of French Roast and one kilogram of Boxer Blend from the Java Hut. Jacob buys two kilograms of French Roast and three kilograms of Boxer Blend. If Joanne spends $14 for her purchase, and Jacob spends $36 for his, what is the price of one kilogram of Boxer Blend coffee?

OpenStudy (anonymous):

x : french y : boxer joanne: x + y = 14 jacob: 2x + 3y = 36 i would suggest rearranging joanne's to x = 14 - y, then plugging that in for x in jacob's equation. this will give you the value of y to plug in either one after.

OpenStudy (anonymous):

x = 7 and y = 7

OpenStudy (anonymous):

I need the price, im just having lots of troubles.

OpenStudy (anonymous):

y = 1kg of boxer the answer in units is the answer in dollars

OpenStudy (anonymous):

only one way to learn is to do it :) jacob: 2(14-y) + 3y = 36. i know you can do it. made a mistake in original

OpenStudy (anonymous):

I have 2y-28+3y=36?

OpenStudy (anonymous):

i made a mistake so its 28-2y+3y=36. you were right but i made a mistake. so if u move stuff around you get y = 8

OpenStudy (anonymous):

i inverted y and 14 the 2nd time i wrote it